A solution is prepared by dissolving 104.1 grams of barium chloride, BaCl2 in 1000. grams of water? What is the freezing point of this solution? [The freezing point depression constant for water is 1.86°C/mole solute in 1000g of water
1
Expert's answer
2019-10-15T07:44:57-0400

"BaCl_2 \to Ba^{+2}+2Cl^-"

t=0 1 0 0

t=T 0 1 2

Vant Hoff's factor for "BaCl_2(i)=\frac{2+1}{1}=3"

Mass of solute"=104.1 g"

Number of moles of solute"=\frac{104.1}{208.2}=0.5" (where 208.2 gm is molar mass of solute)

Mass of solvent"=1000g"

Molality for solution"(m)=\frac{104.1}{1000}\times1000=0.5"

Depression in freezing point"=iK_fm=3\times 1.86 \times 0.5=3 \times 0.93\degree C=2.79\degree C"

So freezing point of water is"\ 0-2.79=-2.79\degree C"